OpenStudy (anonymous):

I WILL MEDAL Select ALL equations that have no solution. A.)6x - 2 - 3x = 3x - 2 B.)6x - (3x + 8) = 16x C.)10 + 6x = 15 + 9x - 3x D.)11 + 3x - 7 = 6x + 5 - 3x

OpenStudy (calculusxy):

No solutions means that the values of x in both sides of the equal sign don't match with each other. For example: 2x + 3 = 2x + 5 2x - 2x + 3 = 2x - 2x + 5 3 = 5 <--- that's not true

OpenStudy (calculusxy):

Just solve for the x's in both sides of the equal sign for each equation. Whichever ones have x's that don't match, then you know that there is no solution.